OpenStudy (anonymous):

if an octahedral die is rolled 9 times, what is the probability that only the last 2 rolls will be a multiple of 3?

OpenStudy (oaktree):

If there are 2 multiples of three out of 8, then it's 2/8 x 2/8 = 1/16.

OpenStudy (anonymous):

hmmm

OpenStudy (oaktree):

But then you need to guarantee that the others are NOT multiples of three, so...

OpenStudy (oaktree):

1/16 x (6/8)^7 = 2187/262144 = approximately 0.008342742919921875

OpenStudy (anonymous):

Thanks for the help OakTree!